# Dual Caption Preference Optimization for Diffusion Models

### v1

@@ Abstract
Recent advancements in human preference optimization, originally developed for Large Language Models (LLMs), have shown significant potential in improving text-to-image diffusion models. These methods aim to learn the distribution of preferred samples while distinguishing them from less preferred ones. However, existing preference datasets often exhibit overlap between these distributions, leading to a conflict distribution. Additionally, we identified that input prompts contain irrelevant information for less preferred images, limiting the denoising network’s ability to accurately predict noise in preference optimization methods, known as the irrelevant prompt issue. To address these challenges, we propose Dual Caption Preference Optimization (DCPO), a novel approach that utilizes two distinct captions to mitigate irrelevant prompts. To tackle conflict distribution, we introduce the Pick-Double Caption dataset, a modified version of Pick-a-Pic v2 with separate captions for preferred and less preferred images. We further propose three different strategies for generating distinct captions: captioning, perturbation, and hybrid methods. Our experiments show that DCPO significantly improves image quality and relevance to prompts, outperforming Stable Diffusion (SD) 2.1, $\text{SFT}_{\text{Chosen}}$, Diffusion-DPO and MaPO across multiple metrics, including Pickscore, HPSv2.1, GenEval, CLIPscore, and ImageReward, fine-tuned on SD 2.1 as the backbone. Our code and dataset are available in Github.
@@ 1 Introduction
Figure 2: The overview of our 3 Dual Preference Optimization (DCPO) pipelines: DCPO-c, DCPO-p, and DCPO-h, all of which require a duo of a captioned preferred image $(x^{w}_{0},z^{w})$ and a captioned less-preferred image $(x^{l}_{0},z^{l})$. DCPO-c (Top Left): We use a captioning model to generate distinctive captions respectively for images $x^{w}_{0}$ and $x^{l}_{0}$ given the shared prompt $c$. DCPO-p (Bottom Left): We take prompt $c$ as the caption for image $x^{w}_{0}$, then we use a Large Language Model (LLM) to generate a semantically perturbed prompt $z_{p}^{l}$ given prompt $c$ as the caption for image $x^{l}_{0}$. DCPO-h (Right): A hybrid method where the generated caption $z^{l}$ is now perturbed into $z_{p}^{l}$ for image $x^{l}_{0}$. Our Pick-Double Caption Dataset discussed in Section 3.1 is constructed with the DCPO-c pipeline.
@@
Image synthesis models (Rombach et al., 2022; Esser et al., 2024) have achieved remarkable advancements in generating photo-realistic and high-quality images. Text-conditioned diffusion (Song et al., 2020a) models have led this progress due to their strong generalization abilities and proficiency in modeling high-dimensional data distributions. As a result, they have found wide range of applications in image editing (Brooks et al., 2023), video generation (Wu et al., 2023a) and robotics (Carvalho et al., 2023). Consequently, efforts have focused on aligning them with human preferences, targeting specific attributes like safety (Liu et al., 2024b), style (Patel et al., 2024; Everaert et al., 2023), spatial understanding (Chatterjee et al., 2024b), and personalization (Ruiz et al., 2023), thereby improving their usability and adaptability.
@@
Similar to the alignment process of Large Language Models (LLMs), aligning diffusion models involves two main steps: 1. Pre-training and 2. Supervised Fine-Tuning (SFT). Recent fine-tuning based methods have been introduced to optimize diffusion models according to human preferences by leveraging Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), the aim of which is to maximize an explicit reward. However, challenges such as reward hacking have led to the adoption of Direct Preference Optimization (DPO) (Rafailov et al., 2024) techniques like Diffusion-DPO (Wallace et al., 2024). Intuitively, Diffusion-DPO involves maximizing the difference between a preferred and a less preferred image for a given prompt.
@@
Although DPO-based methods are effective in comparison to SFT-based approaches, applying direct optimization in multi-modal settings presents certain challenges. Current preference optimization datasets consist of a preferred ($x^{w}$) and a less preferred ($x^{l}$) image for a given prompt ($c$). Ideally, $x^{w}$ should show a higher correlation with $c$ compared to $x^{l}$. However, we find that in current datasets, both the images share the same distribution for the given prompt $c$, which we refer to as conflict distribution in the data. Additionally, irrelative information in $c$ restricts the U-Net’s ability to predict noises from $x^{l}$ in the diffusion reverse process, which we refer to as irrelevant prompts. This entails that there is a lack of sufficient distinguishing features between the two pairs ($x^{w}$, $c$), ($x^{l}$, $c$), thereby increasing the complexity of the optimization process.
Overall, we identify two key challenges: 1. conflict distribution in data, which contradicts the core idea of direct optimization, and 2. irrelevant prompts for less preferred images, which can hinder the learning process of the diffusion model during optimization.
@@
To address the aforementioned bottleneck, we propose DCPO: Dual Caption Preference Optimization, a novel preference optimization technique designed to align diffusion models by utilizing two distinct captions corresponding to the preferred and less preferred image. DCPO broadly consists of two steps - a text generation framework that develops better aligned captions and a novel objective function that utilizes these captions as part of the training process.
@@
The text generation framework seeks to alleviate the conflict distribution present in existing datasets. We hypothesize that $c$ does not serve as the optimal signal for optimization because they do not convey the reasons why an image is preferred or dis-preferred; based on the above, we devise the following techniques to generate better aligned captions. The first method involves using a captioning model $Q_{\phi}(z^{i}|x^{i},c)$; which generates a new prompt ${z^{i}}$ based on an image ${x^{i}}$ and the original prompt ${c}$, where $i\in(w,l)$. The second method introduces perturbation techniques $f$, such that $c=z^{w},z^{l}=f(c);$ i.e. generating $z^{l}$, to represent the less preferred image, considering the original prompt $c$ as the prompt aligned with the preferred image. We investigate multiple semantic variants of $f$, where each variant differs in the degree of perturbation applied to the original caption $c$. Finally, we also explore a hybrid combination of the above methods, where we combine the strong prior of the captioning model and the efficient nature of the perturbation method. All the above methods are designed to generate captions that effectively discriminate between the preferred and less preferred images.
@@
We introduce a novel objective function that allows DCPO to incorporate $z^{w}$ and $z^{l}$ into its optimization process. Specifically, during optimization, the policy model $p_{\theta}$ increases the likelihood of the preferred image $x^{w}$ conditioned on the prompt $z^{w}$, while simultaneously decreasing the likelihood of the less preferred image $x^{l}$ conditioned on the prompt $z^{l}$. The results in Tables 1 and 2 demonstrate that DCPO consistently outperforms other methods, with notable improvements of +0.21 in Pickscore, +0.45 in HPSv2.1, +1.8 in normalized ImageReward, +0.15 in CLIPscore, and +3% in GenEval. Additionally, DCPO achieved 58% in general preference and 66% in visual appeal compared to Diffusion-DPO on the PartiPrompts dataset, as evaluated by GPT-4o (see Figure 8).
@@
  • Double Caption Generation: We introduce the Captioning and Perturbation methods to address the conflict distribution issue, as illustrated in Figure 3. In the Captioning method, we employ state-of-the-art models like LLaVA (Liu et al., 2024a) and Emu2 (Sun et al., 2024) to generate a caption $z$ based on the image $x$ and prompt $c$. Additionally, we use DIPPER (Krishna et al., 2024), a paraphrase generation model built by fine-tuning the T5-XXL model to create three levels of perturbation from the prompt $c$.
@@ 2 Method
In this section, we present the conflict distribution issue in preference datasets, where preferred and less-preferred images generated from the same prompt $c$ exhibit significant overlap. We also explain the irrelevant prompt issue found in previous direct preference optimization methods. To address these challenges, we propose Dual Caption Preference Optimization (DCPO), a method that uses distinct captions for preferred and less preferred images to improve diffusion model alignment.
@@ 2.1 The Challenges
Figure 3: The conflict distribution issue in the Pick-a-Pic v2 dataset. $\mu^{l}$ and $\mu^{w}$ represent the average CLIPscore of preferred and less preferred images for prompt $c$, respectively. Also, $\Delta\mu$ shows the difference between the distributions.
Generally, to optimize a Large Language Model (LLM) using preference algorithms, we need a dataset $D=\{c,y^{w},y^{l}\}$, where $y^{w}$ and $y^{l}$ represent the preferred and less preferred responses to a given prompt $c$. Ideally, the distributions of these responses should differ significantly. Similarly, in diffusion model alignment, the distributions of preferred and less preferred images should be distinct for the same prompt $c$. However, our analysis shows a substantial overlap between these distributions, which we call conflict distribution, as illustrated in Figure 3.
@@ 2.2 DCPO: Dual Caption Preference Optimization
Motivated by the conflict distribution and irrelevant prompts issues, we propose DCPO, a new preference optimization method that optimizes diffusion models using two distinct captions. DCPO is a refined version of Diffusion-DPO designed to address these challenges. More details are in Appendix B.  [^1] For additional background about diffusion and preference optimization, refer to Appendix A.
@@
$\begin{split}\mathcal{L}_{\text{DCPO}}(\theta)=-\mathit{\mathbb{E}}_{(x^{w}_{0},x^{l}_{0},z^{l},z^{w})\sim\mathcal{D^{\prime}}}\log\sigma(\beta\mathit{\mathbb{E}}_{x^{w}_{1:T}\sim p_{\theta}(x^{w}_{1:T}|x^{w}_{0},z^{w}),x^{l}_{1:T}\sim p_{\theta}(x^{l}_{1:T},x^{l}_{0}|z^{l})}\\ [\log\frac{p_{\theta}(x^{w}_{0:T}|z^{w})}{p_{\text{ref}}(x^{w}_{0:T}|z^{w})}-\log\frac{p_{\theta}(x^{l}_{0:T}|z^{l})}{p_{\text{ref}}(x^{l}_{0:T}|z^{l})}])\end{split}$ (1)
However, as noted in Diffusion-DPO (Wallace et al., 2024), the sampling process $x_{1:T}\sim p(x_{1:T}|x_{0})$ is inefficient and intractable. To overcome this, we follow a similar approach by applying Jensen’s inequality and utilizing the convexity of the $-\log(\cdot)$ function to bring the expectation outside. By approximating the reverse process $p_{\theta}(x_{1:T}|x_{0},z)$ with the forward process $q(x_{1:T}|x_{0})$, and through algebraic manipulation and simplification, the DCPO loss can be expressed as:
$\begin{split}\mathcal{L}_{\text{DCPO}}(\theta)=-\mathit{\mathbb{E}}_{(x^{w}_{0},x^{l}_{0})\sim\mathcal{D^{\prime}},t\sim\mu(0,T),x^{w}_{t}\sim q(x^{w}_{t}|x^{w}_{0}),x^{l}_{t}\sim q(x^{l}_{t}|x^{l}_{0})}\\ \log\sigma(-\beta Tw(\lambda_{t})(||\epsilon^{w}-\epsilon_{\theta}(x^{w}_{t},z^{w},t)||^{2}_{2}-||\epsilon^{w}-\epsilon_{\text{ref}}(x^{w}_{t},z^{w},t)||^{2}_{2}\\ -(||\epsilon^{l}-\epsilon_{\theta}(x^{l}_{t},z^{l},t)||^{2}_{2}-||\epsilon^{l}-\epsilon_{\text{ref}}(x^{l}_{t},z^{l},t)||^{2}_{2}))\end{split}$ (2)
@@ 2.2.1 DCPO-c: Captioning Method
In the experiment section, we evaluate the performance of DCPO-c and demonstrate that this method effectively mitigates the conflict distribution by creating two differentiable distributions. However, the question of how much divergence is needed between the two distributions remains. To investigate this, we propose Hypothesis 1.
Hyphothesis 1. Let $d(z,x)$ represent the semantic distribution between a caption $z$ and an image $x$, with $\mu$ being the mean of the distribution $d$, and $\Delta\mu=\mu(d(z_{0}^{w},x_{0}^{w}))-\mu(d(z_{0}^{l},x_{0}^{l}))$ as the difference between the two distributions. Increasing $\Delta\mu$ between the preferred and less-preferred image distributions in a preference dataset beyond a threshold $t$ (i.e., $\Delta\mu>t$), can improve the performance of the model $p_{\theta}$.
@@ 2.2.2 DCPO-p: Perturbation Method
Figure 4: Effect of the perturbation method on semantic distributions in terms of CLIPScore. (a) shows the distributions that feature the captions $z^{w}$ and $z^{l}$ generated by the LLaVA model, while (b), (c), and (d) represent different levels of perturbation on top of the caption $z^{l}$. The figure demonstrates that as the level of perturbation increases, the distance between the distributions of captions $z^{w}$ and $z^{l}$ increases. For more details on the perturbation method, refer to Appendix E.
@@
While using a captioning model is an effective way to address the conflict distribution, it risks deviating from the original distribution of prompt $c$, and the distributions of preferred and less preferred images may still remain close. To tackle these issues, we propose a perturbation method. In this approach, we assume that prompt $c$ is highly relevant to the preferred image $x_{0}^{w}$ and aim to generate a less relevant caption, denoted as $c_{p}$, based on prompt $c$. To achieve this, we use the model $W_{\phi}(c_{p}|c)$, which generates a perturbed version of prompt $c$, altering its semantic meaning. In this framework, prompt $c$ corresponds to the preferred caption $z^{w}$ ($c=z^{w}$), while the perturbed prompt $c_{p}$ represents the less-preferred caption $z^{l}$ ($c_{p}=z^{l}$). For the perturbation model $W_{\phi}$, we utilized the DIPPER model (Krishna et al., 2024) built by fine-tuning the T5-XXL  (Chung et al., 2022) to produce a degraded version of the prompt $c$.
@@
We define three levels of perturbation: 1) Weak: where prompt $c_{p}$ has high semantic similarity to prompt $c$, with minimal differences. 2) Medium: where the semantic difference between prompt $c_{p}$ and $c$ is more pronounced than in the weak level. 3) Strong: where the majority of the semantics in prompt $c_{p}$ differ significantly from prompt $c$. Further details can be found in Appendix E.
@@
The main advantage of DCPO-p is to reduce the captioning process cost while staying closer to the original data distribution by using prompt $c$ as the preferred caption. However, we observe that the quality of captions in DCPO-c outperforms that of the original prompt $c$, as shown in Table 6 in Appendix D. Based on this observation, we propose a hybrid method to improve the alignment performance by combining captioning and perturbation techniques.

### v2

@@ Abstract
Recent advancements in human preference optimization, originally developed for Large Language Models (LLMs), have shown significant potential in improving text-to-image diffusion models. These methods aim to learn the distribution of preferred samples while distinguishing them from less preferred ones. However, within the existing preference datasets, the original caption often does not clearly favor the preferred image over the alternative, which weakens the supervision signal available during training. To address this issue, we introduce Dual Caption Preference Optimization (DCPO), a data augmentation and optimization framework that reinforces the learning signal by assigning two distinct captions to each preference pair. This encourages the model to better differentiate between preferred and less-preferred outcomes during training. We also construct Pick-Double Caption, a modified version of Pick-a-Pic v2 with separate captions for each image, and propose three different strategies for generating distinct captions: captioning, perturbation, and hybrid methods. Our experiments show that DCPO significantly improves image quality and relevance to prompts, outperforming Stable Diffusion (SD) 2.1, $\text{SFT}_{\text{Chosen}}$, Diffusion-DPO, and MaPO across multiple metrics, including Pickscore, HPSv2.1, GenEval, CLIPscore, and ImageReward, fine-tuned on SD 2.1 as the backbone. Our code and dataset are available in Github.
@@ 1 Introduction
Figure 2: The DCPO pipeline in 3 variants: DCPO-c, DCPO-p, and DCPO-h, all of which require a duo of a captioned preferred image $(x^{w}_{0},z^{w})$ and a captioned less-preferred image $(x^{l}_{0},z^{l})$. DCPO-c (Top Left): We use a captioning model to generate distinctive captions respectively for images $x^{w}_{0}$ and $x^{l}_{0}$ given the shared prompt $c$. DCPO-p (Bottom Left): We take prompt $c$ as the caption for image $x^{w}_{0}$, then we use a Large Language Model (LLM) to generate a semantically perturbed prompt $z_{p}^{l}$ given prompt $c$ as the caption for image $x^{l}_{0}$. DCPO-h (Right): A hybrid method where the generated caption $z^{l}$ is now perturbed into $z_{p}^{l}$ for image $x^{l}_{0}$. Our Pick-Double Caption Dataset discussed in Section 4.1 is constructed using DCPO-c.
@@
Image synthesis models (Rombach et al., 2022; Esser et al., 2024) have achieved remarkable advancements in generating photo-realistic and high-quality images. Text-conditioned diffusion (Song et al., 2020) models have led this progress due to their strong generalization abilities and proficiency in modeling high-dimensional data distributions. As a result, they have found wide range of applications in image editing (Brooks et al., 2023), video generation (Wu et al., 2023a) and robotics (Carvalho et al., 2023). Consequently, efforts have focused on aligning them with human preferences, targeting specific attributes like safety (Liu et al., 2024b), style (Everaert et al., 2023), and personalization (Ruiz et al., 2023), thereby improving their usability and adaptability.
@@
Similar to the alignment process of Large Language Models (LLMs), aligning diffusion models involves two main steps: 1. Pre-training and 2. Supervised Fine-Tuning (SFT). Recent fine-tuning based methods have been introduced to optimize diffusion models according to human preferences by leveraging Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), the aim of which is to maximize an explicit reward. However, challenges such as fine-tuning a separate reward model and reward hacking have led to the adoption of Direct Preference Optimization (DPO) (Rafailov et al., 2024) techniques like Diffusion-DPO (Wallace et al., 2024). Intuitively, Diffusion-DPO involves maximizing the difference between a preferred image and a less preferred image for a given prompt.
@@
Although DPO-based methods are effective in comparison to SFT-based approaches, applying direct optimization in multi-modal settings presents certain challenges. Current preference optimization datasets consist of a preferred ($x^{w}$) and a less preferred ($x^{l}$) image for a given prompt ($c$). Ideally, $x^{w}$ should show a higher correlation with $c$ compared to $x^{l}$. However, we find that in current datasets, the preferred and less-preferred images exhibit a substantial overlap in their semantic distributions for the given prompt $c$, which we refer to as semantic overlap in the data. Additionally, irrelevant information in prompt $c$ restricts the U-Net’s ability to predict noises from $x^{l}$ in the diffusion reverse process, which we refer to as irrelevant prompts. This entails that there is a lack of sufficient distinguishing features between the two pairs ($x^{w}$, $c$), ($x^{l}$, $c$), thereby increasing the complexity of the optimization process.
@@
To address the aforementioned bottleneck, we propose DCPO: Dual Caption Preference Optimization, a novel preference optimization technique designed to align diffusion models by utilizing two distinct captions corresponding to the preferred and less preferred image. DCPO broadly consists of two steps: a text generation framework that develops better-aligned captions and a novel objective function that utilizes these captions as part of the training process.
@@
The text generation framework aims to mitigate the semantic overlap between the distributions of preferred and less-preferred images, given the same prompts, in existing datasets. We hypothesize that prompt $c$ does not serve as the optimal signal for optimization because they do not convey the reasons why an image is preferred or dis-preferred; based on the above, we devise the following techniques to generate better aligned captions. The first method involves using a captioning model $Q_{\phi}(z^{i}|x^{i},c)$; which generates a new prompt $z^{i}$} based on an image $x^{i}$} and the original prompt c{c}, where $i\in(w,l)$. The second method introduces perturbation techniques $f$, such that $c=z^{w},z^{l}=f(c);$ i.e., generating $z^{l}$, to represent the less preferred image, considering the original prompt $c$ as the prompt aligned with the preferred image. We investigate multiple semantic variants of $f$, where each variant differs in the degree of perturbation applied to the original caption $c$. Finally, we also explore a hybrid combination of the above methods, where we combine the strong prior of the captioning model and the efficient nature of the perturbation method. All the above methods are designed to generate captions that effectively discriminate between the preferred and less preferred images.
@@
We introduce a novel objective function that allows DCPO to incorporate $z^{w}$ and $z^{l}$ into its optimization process. Specifically, during optimization, the policy model $p_{\theta}$ increases the likelihood of the preferred image $x^{w}$ conditioned on the prompt $z^{w}$, while simultaneously decreasing the likelihood of the less preferred image $x^{l}$ conditioned on the prompt $z^{l}$. The results in Tables 1 and 2 demonstrate that DCPO consistently outperforms other methods, with notable improvements of +0.21 in Pickscore, +0.31 in HPSv2.1, +1.8 in normalized ImageReward, +0.15 in CLIPscore, and +2% in GenEval. Additionally, DCPO achieved 58% in general preference and 66% in visual appeal compared to Diffusion-DPO on the PartiPrompts dataset, as evaluated by GPT-4o (see Figure 7).
@@
  • Double Caption Generation: We introduce the Captioning and Perturbation methods to address the issue of overlapping semantic distributions in current datasets, as illustrated in Figure 3. In the Captioning method, we employ state-of-the-art models like LLaVA (Liu et al., 2024a) and Emu2 (Sun et al., 2024) to generate a caption $z$ based on the image $x$ and prompt $c$. Additionally, we use DIPPER (Krishna et al., 2024), a paraphrase generation model built by fine-tuning the T5-XXL model to create three levels of perturbation from the prompt $c$.
@@
2 Preliminary
$p_{BT}(x^{w}_{0}\succ x^{l}_{0}|c)=\sigma(r(c,x^{w}_{0})-r(c,x^{l}_{0}))$ (1)
where $\sigma(\cdot)$ is the sigmoid function, and $r(x_{0},c)$ is derived from a neural network parameterized by $\phi$.
Subsequently, $\phi$ is estimated by maximum likelihood training for binary classification as follows:
where prompt $c$ and data pair $(x_{0}^{w},x_{0}^{l})$ are sourced from a human-annotated dataset.
where $\beta$ controls how far the policy model $p_{\theta}$ can deviate from the reference model $p_{ref}$.
It can be demonstrated that the objective in Equation 3 converges to the following policy model:
where $Z$ is the partition function.
The training objective for $p_{\theta}$, inspired by DPO, has been derived to be equivalent to Equation 4 without the need for an explicit reward model $r(x,c)$. Instead, RLHF learns directly from the preference data $(c,x^{w}_{0},x^{l}_{0})\sim D$:
$L_{\text{DPO}}(\theta)=-\mathit{\mathbb{E}}_{c,x^{w}_{0},x^{l}_{0}}\Big[\log\sigma\Big(\beta\log\frac{p_{\theta}(x^{w}_{0}|c)}{p_{\text{ref}}(x^{w}_{0}|c)}-\beta\log\frac{p_{\theta}(x^{l}_{0}|c)}{p_{\text{ref}}(x^{l}_{0}|c)}\Big)\Big]$ (5)
where $\sigma(\cdot)$ is the sigmoid function.
Through this re-parameterization, instead of optimizing the reward function $r$ before applying reinforcement learning, the RLHF method directly optimizes the conditional distribution $p_{\theta}(x_{0}|c)$.
@@ 3 Method
In this section, we present the semantic overlap issue preference datasets, where the semantic distribution of preferred and less-preferred images generated from the same prompt $c$ exhibit significant similarity. We also explain the irrelevant prompt issue found in previous direct preference optimization methods. To address these challenges, we propose Dual Caption Preference Optimization (DCPO), a method that uses distinct captions for preferred and less preferred images to improve diffusion model alignment.
@@ 3.1 The Challenges
Generally, to optimize a Large Language Model (LLM) using preference algorithms, we need a dataset $D=\{c,y^{w},y^{l}\}$, where $y^{w}$ and $y^{l}$ represent the preferred and less preferred responses to a given prompt $c$. Recent studies (Sun et al., 2025; Deng et al., 2025) have shown that high semantic similarity between preferred and less-preferred responses negatively impacts preference optimization and should be carefully considered during dataset curation. Similarly, in diffusion model alignment, the semantic distributions of preferred and less preferred images should be distinct for the same prompt $c$. However, our analysis shows a substantial overlap between these distributions, which we call semantic overlap, as illustrated in Figure 3. For more details, refer to Appendix B.1.
Figure 3: The semantic overlap issue in the Pick-a-Pic v2 dataset. $\mu^{l}$ and $\mu^{w}$ represent the average CLIPscore of preferred and less preferred images for prompt $c$, respectively. Also, $\Delta\mu$ shows the difference between the distributions.
@@ 3.2 DCPO: Dual Caption Preference Optimization
Motivated by the semantic overlap and irrelevant prompts issues, we propose DCPO, a new preference optimization method that optimizes diffusion models using two distinct captions. DCPO is a refined version of Diffusion-DPO designed to address these challenges. More details are in Appendix A.
@@
$\mathcal{L}_{\text{DCPO}}(\theta)=-\mathit{\mathbb{E}}_{(x^{w}_{0},x^{l}_{0})\sim\mathcal{D^{\prime}}}\log\sigma\Big[\beta\mathit{\mathbb{E}}_{x^{w}_{1:T}\sim p_{\theta}(x^{w}_{1:T}|x^{w}_{0},z^{w}),x^{l}_{1:T}\sim p_{\theta}(x^{l}_{1:T}|x^{l}_{0},z^{l})}\Big(\log\frac{p_{\theta}(x^{w}_{0:T}|z^{w})}{p_{\text{ref}}(x^{w}_{0:T}|z^{w})}-\log\frac{p_{\theta}(x^{l}_{0:T}|z^{l})}{p_{\text{ref}}(x^{l}_{0:T}|z^{l})}\Big)\Big]$ (6)
where $\log[\cdot]$ is the sigmoid function.
However, as noted in Diffusion-DPO (Wallace et al., 2024), the sampling process $x_{1:T}\sim p(x_{1:T}|x_{0})$ is inefficient and intractable. To overcome this, we follow a similar approach by applying Jensen’s inequality and utilizing the convexity of the $\log(\cdot)$ function to bring the expectation outside. By approximating the reverse process $p_{\theta}(x_{1:T}|x_{0},z)$ with the forward process $q(x_{1:T}|x_{0})$, and through algebraic manipulation and simplification, the DCPO loss can be expressed as:
[…]
@@ 3.2.1 DCPO-c: Captioning Method
In the experiment section, we evaluate the performance of DCPO-c and demonstrate that this method effectively mitigates the semantic overlap by creating two differentiable distributions. However, the question of how much divergence is needed between the two distributions remains. To investigate this, we propose Hypothesis 1.
Hyphothesis 1.
Let $d(z,x)$ represent the semantic distribution between a caption $z$ and an image $x$, with $\mu$ being the mean of the distribution $d$, and $\Delta\mu=\mu(d(z_{0}^{w},x_{0}^{w}))-\mu(d(z_{0}^{l},x_{0}^{l}))$ as the difference between the two distributions. Increasing $\Delta\mu$ between the preferred and less-preferred image distributions in a preference dataset beyond a threshold $t$ (i.e., $\Delta\mu>t$), can improve the performance of the model $p_{\theta}$.
@@ 3.2.2 DCPO-p: Perturbation Method
Figure 4: Effect of the perturbation method on semantic distributions in terms of CLIPScore. (a) shows the distributions that feature the captions $z^{w}$ and $z^{l}$ generated by the LLaVA model, while (b), (c), and (d) represent different levels of perturbation on top of the caption $z^{l}$. The figure demonstrates that as the level of perturbation increases, the distance between the distributions of captions $z^{w}$ and $z^{l}$ increases. For more details on the perturbation method, refer to Appendix D.
@@
While using a captioning model is an effective way to address the semantic overlap, it risks deviating from the original distribution of prompt $c$, and the distributions of preferred and less preferred images may still remain close. To tackle these issues, we propose a perturbation method. In this approach, we assume that prompt $c$ is highly relevant to the preferred image $x_{0}^{w}$ and aim to generate a less relevant caption, denoted as $c_{p}$, based on prompt $c$. To achieve this, we use the model $W_{\phi}(c_{p}|c)$, which generates a perturbed version of prompt $c$, altering its semantic meaning. In this framework, prompt $c$ corresponds to the preferred caption $z^{w}$ ($c=z^{w}$), while the perturbed prompt $c_{p}$ represents the less-preferred caption $z^{l}$ ($c_{p}=z^{l}$).
For the perturbation model $W_{\phi}$, we choose the highly customizeable DIPPER model (Krishna et al., 2024), which is built by fine-tuning T5-XXL (Chung et al., 2022), to produce a degraded version of the prompt $c$. DIPPER is equipped with flexible steerability for controllable textual perturbation, which supports adjustable hyperparameters that control both lexical diversity and word order variation at the output.
@@
We define three levels of perturbation strength: 1) Weak: where prompt $c_{p}$ has high semantic similarity to prompt $c$, with minimal differences. 2) Medium: where the semantic difference between prompt $c_{p}$ and $c$ is more pronounced than in the weak level. 3) Strong: where the majority of the semantics in prompt $c_{p}$ differ significantly from prompt $c$.
Perturbation Selection Process.
Given each input prompt $c$, we first perform a hyperparameter sweep over DIPPER’s controllable dimensions to construct a larger diverse pool of perturbed prompt candidates. We then select the candidates that are well-separated along the CLIPScore scale from the pool, ensuring a clear semantic gradient while always preserving the original input as the one with the highest image-text alignment in terms of CLIPScore affinity. The three resulting perturbed prompts $c_{p}$ — in levels of Weak, Medium, and Strong, respectively — reflect progressively degraded semantic fidelity relative to the associated image of the input, visualized as the shifting distributions in terms of CLIPScore semantic similarity in Figure 4. For further details on the perturbation method, see Appendix D.
@@
Compared with DCPO-c, DCPO-p is, first of all, less computationally costly. As described in Figure 2, DCPO-p requires only the textual perturbation as the less-preferred caption. In addition, by using the original prompt $c$ directly as the preferred caption, DCPO-p keeps the original distribution at bay, mitigating DCPO-c’s risk where both the preferred and the less preferred may deviate together. However, we observe that the quality of the captions in DCPO-c outperforms that of the original prompt $c$, as shown in Table LABEL:tab:tokens-detail in Appendix C. Based on this observation, we propose a hybrid method to improve the alignment performance by combining the captioning and perturbation techniques from DCPO-c and DCPO-p, altogether.